OpenStudy (anonymous):

Find the two positive numbers whose sum is 110 and whose product is a maximum.

OpenStudy (anonymous):

A maximum of what?

OpenStudy (anonymous):

y + x = 110 = maximum

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

x+y=110; xy=M xy is what you wanna maximize. Solve the first for y y=110-x Plug in your max. (110-x)(x)=M 110x-x^2=M Differentiate: 110-2x=dM/dx Set equal to zero. 110-2x=0 2x=110 x=55 Try pluggin in x=54 and x=56 to make sure that it is a max (First derivative test)